Find the slope of the line whose equation is 3x-2y=8.
Find the equation of a line given that the slope m=-2 and it passes through the point(-2,1).
Find the slope of the line passing through the points(6,5)and (-2,-1).
solve for x:1/x+2/x-2=4/x

Answer by nyc_function(2733)

BIG TIP: POST ONE QUESTION AT A TIME.
Find the slope of the line whose equation is 3x-2y=8.
3x - 2y = 8
Our goal is to isolate y. The slope will be the number next to x.
-2y = -3x + 8
y = (-3x + 8)/-2
y = (3/2)(x) - 4
The slope is 3/2